A box of bulbs is inspected. The inspector draws a bulb randomly and independently, and tests it until he finds 10 defective bulbs. Suppose 5% of all bulbs are defective. Let X be the number of bulbs tested in order to find 10 defective bulbs.

• A) What is the distribution of X?

• B) What is the expected number of bulbs that must be tested in order to get 10 defective bulbs?

• C) What is the variance of X?

• D) What is the probability that 15 bulbs will be tested in order to find

10 defective bulbs?



1
Expert's answer
2022-05-30T09:34:28-0400

A) A negative binomial experiment is a statistical experiment that has the following properties:

The experiment consists of x repeated trials.

Each trial can result in just two possible outcomes. We call one of these outcomes a success and the other, a failure.

The probability of success, denoted by "p," is the same on every trial.

The trials are independent; that is, the outcome on one trial does not affect the outcome on other trials.

The experiment continues until "r" successes are observed, where "r"  is specified in advance.

"X\sim NB(10 ,0.05)"

B)


"E(X)=\dfrac{r(1-p)}{p}=\dfrac{10(1-0.05)}{0.05}=190"

C)


"Var(X)=\dfrac{r(1-p)}{p^2}=\dfrac{10(1-0.05)}{(0.05)^2}=3800"

D)


"nb(10;5,0.05)=\dbinom{15-1}{10-1}(0.05)^{10}(1-0.05)^{15-10}"

"=1.5\times10^{-10}"